# Theory of the Spin Seebeck effect induced by crystal-field excitations in Tb3Fe5O12

### v2

@@ Appendix B Magnetic field dependence of the crystal-field excitations
Fig. 5 is consistent with Fig. 8 for $B_{111}>14.4$ T. This is because the applied field $B$ of Fig. 8  is added to the molecular field $B^{z}_{mf}$ to give the total field, both in the [111] directions, for this figure.
Figure 8: The magnetic field dependence of the lowest six CF excitations of the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites under a total magnetic field applied along the [111] direction. Fig.  5 in the main text corresponds to the detail of this figure, with the origin of the horizontal-axis shifted by our estimated molecular field $14.4$ T; see Eq.23).The CF parameters used are from TGG; see left column of Table 1 [21].
To verify the robustness of the magnetic field dependences, we studied several directions of the applied fields, also fairly off from the [111] direction. Motivated by the enhanced SSE reported in Ref. 16 for TbIG at high temperatures and fields applied along the crystallographic [100], we show in Fig. 9 the single-ion results from exact diagonalisation of the Hamiltonian
$H_{CF}-g_{J}\mu_{\mathrm{B}}\,\hat{\mathbf{J}}\cdot\left(B_{0}\,\frac{[111]}{\sqrt{3}}+B\,[100]\right),$ (24)
with $B_{0}=15$ T, fixed to incorporate the effect from the [111] molecular field of $\simeq$14.414.414.414.4 T, as in Eq.23, and $B$ the tunable strength of the [100] applied field (see horizontal axis of Fig. 9). Our results also in this case corroborate the anomalous behaviour of the lowest CFE (red curve), whose gap with the ground state decreases over a range of more than 5 T for sites 1, 2, and 5 (see top panel of Fig. 9), whilst it only increases for about 10 T for the sites 3 and 6 (see mid-panel of Fig. 9) and as well for site 4 (see bottom panel of Fig. 9).
Figure 9: The magnetic field dependence of the lowest three CF excitations from exact diagonalisation of Eq. (24). The molecular field from the Fe-sites is fixed along the [111] direction ($B_{0}=15$ T); in addition, we have the externally applied field $B$ along the [100] direction. Note in comparing Figs. 5, 8 and 9 that energies on the axis $B=0$ in Figs. 5 and 9 correspond to those of $B=14.4$ T in Fig. 8. The CF parameters used are from TGG; see left column of Table 1 [21].

### v3

@@ Appendix C Magnetic field dependence of the crystal-field excitations
Fig. 5 is consistent with Fig. 9 for $B\,[111]>14.4$ T. This is because the applied field $B$ of Fig. 9  is added to the molecular field $B^{z}_{mf}$ to give the total field, both in the [111] directions, for this figure.
Figure 9: The magnetic field dependence of the lowest five CF excitations of the $\mathbf{C}$ and $\mathbf{C^{\prime}}$ sites under a total magnetic field applied along the [111] direction. Fig.  5 in the main text corresponds to the detail of this figure, with the origin of the horizontal-axis shifted by our estimated molecular field $14.4$ T; see Eq. (23). The broken line with $B\,[111]=15$ T is put as a rough estimate of the molecular field 14.4 T, and will be referred in Fig. 10. The CF parameters used are from TGG; see left column of Table 1 [24].
To explore the magnetic field dependency, we studied several directions of the applied fields, also departing from the [111] direction, but assuming that the molecular field remains in the [111]. In particular, motivated by the enhanced SSE reported in Ref. 19 for TbIG at high temperatures and fields applied along the crystallographic [100], we show in Fig. 10 the single-ion results from the Hamiltonian
$H_{\rm[100]}=H_{CF}-g_{J}\mu_{\mathrm{B}}\,\hat{\mathbf{J}}\cdot\Bigl{(}\mathbf{B}\,[111]+\mathbf{B}\,[100]\Bigr{)},$ (24)
with $B\,[111]=15$ T, fixed to incorporate the effect from the [111] molecular field of $\simeq$14.414.414.414.4 T, as in Eq. 23, and $B\,[100]$ the tunable strength of the [100] applied field (see horizontal axis of Fig. 10). In this case the anomalous behaviour of the lowest CFE (red curve), with a gap that decreases over a range of more than 5 T, is only seen for sites 1 and 2, while the lowering of symmetry now gives different gaps for site 3 which are all increasing, (see second panel of Fig. 10). For sites 4 and 5, the lowest gap now decreases, but starting from the higher value at zero external field, (see the third -panel of Fig. 10) and for site 6 (see bottom panel of Fig. 10) it increases.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 10: The magnetic field dependence of the CF excitations from Eq. (24). In each figure, the field dependence with $B\,[111]$ in Fig. 9 is also shown to see how each curve changes. The molecular field from the Fe-sites along the [111] direction is fixed to $B\,[111]=15$ T at the solid line; in addition, we have the externally applied field $B$ along the [100] direction, $B\,[100]$. The CF parameters used are from TGG; see left column of Table 1 [24].